OpenStudy (anonymous):

derivative of cos(2t) is?

OpenStudy (anonymous):

-2sin(2t)

OpenStudy (anonymous):

t doesnt come out, huh. thanx!

OpenStudy (anonymous):

would 3cos(2t) be 6cos(2t)?

OpenStudy (anonymous):

i mean -6sin(2t)

OpenStudy (amistre64):

the t DOES come out, but under disguise

OpenStudy (amistre64):

D[cos(2t)] = -cos(2t) * D[2t] D[cos(2t)] = -sin(2t) * 2 * D[t] D[cos(2t)] = -sin(2t) * 2 * 1

OpenStudy (anonymous):

what about asin2t-bcos2t i got 2acos2t + 2bsin2t